# A Neural Material Point Method for Particle-based Simulations

### v2

@@ 4 Experiments
We conduct a series of experiments to demonstrate the accuracy, speed, and generalization capabilities of NeuralMPM. Specifically, we examine its robustness to hyperparameter and architectural choices through an ablation study (4.1). We compare NeuralMPM to GNS and DMCF in terms of accuracy, training time, convergence, and inference speed (4.2). We also evaluate the generalization capabilities of NeuralMPM (4.3) and illustrate how its differentiability can be leveraged to solve an inverse design problem (4.4). Through these experiments, we demonstrate that NeuralMPM is a flexible, accurate, and fast method for emulating complex particle-based simulations. The baselines established by Winchenbach & Thuerey (2024) and hybrid simulators (Li et al., 2024, Ma et al., 2023) have promising results. However, we do not compare against them as they either use different benchmarks or are specifically tailored for certain physical domains, requiring material-specific knowledge. In contrast, NeuralMPM, like GNS and DMCF, requires only particle positions without being restricted to any particular domain.
@@
Figure 3: Ablation results. Mean squared error (MSE) of full rollouts on unseen test data for Goop. The default parameters are in blue. The dashed orange line shows the best rollout MSE of GNS as reported by Sanchez-Gonzalez et al. (2020) on the same data ($1.89\times 10^{-3}$), while the dotted line ($2.4\times 10^{-3}$) indicates the MSE we obtained for GNS after $240$ hours ($20$M training steps). NeuralMPM is robust to hyperparameter changes, with the biggest effects coming from the number of timesteps bundled together ($m$) and grid noise. For a rollout of length $T$, the model is called $T/m$ times, meaning lower values of $m$ require maintaining stability for longer. Autoregressive training coupled with time bundling is sufficient to stabilize the model, eliminating the need for noise injection. Although GNS reportedly outperforms NeuralMPM by a small margin, these results could not be reproduced in our experiments.
Figure 4: Example WaterRamps trajectory against baselines. Each method is unrolled starting from the initial conditions of a random test trajectory not seen during training. We display six snapshots evenly spaced in time.
@@ 4.1 Ablation study
To study the robustness of NeuralMPM to hyperparameter and architectural choices, we start with the default architecture and hyperparameters and ablate its components individually to examine their impact on performance. We vary the number $K$ of autoregressive steps with and without noise, the number of bundled timesteps $m$ predicted by a single model call, and the depth and number of hidden channels of the network. We also investigate adding noise to stabilize rollouts, either directly to the particles’ positions or to the grid-level representation after voxelization.
Figure 3 summarizes the ablation results. A larger number $K$ of autoregressive steps yields more accurate rollouts without the need to add noise. Indeed, injecting noise does not improve accuracy and is even detrimental for $K=4$. Individually tuning the noise levels for grids and particles can modestly lower error rates, but is either very sensitive or negligible. The model performs better when bundling more timesteps, enabling faster rollouts as a single forward pass predicts more steps. We found $m=8$ to be optimal with the other default hyperparameters, outperforming larger bundling. This is because more network capacity is needed to extract information for the next $16$ or $32$ timesteps from a single state. Instead, we opted for a shallower and narrower network to balance speed and memory footprint with performance gains. In terms of network architecture, we chose a U-Net. We experiment with an FNO (Li et al., 2021) in Appendix B and find it to underperform, leading us to keep the U-Net architecture. We find the U-Net’s width and depth to have a minor impact on performance, confirming that a larger network is not needed. The grid size, however, is critical. A low resolution loses fine details, while a high resolution turns meaningful structures, such as liquid blobs or walls, into isolated voxels.
@@ 4.2 Comparison with previous work
Figure 6: Time and memory performance. Average FPS (left) and GPU VRAM usage (right) for increasing numbers of particles for a traditional solver (Taichi-MPM (Hu et al., 2018a)), NeuralMPM, and the two baselines. The two baselines quickly require very large amounts of memory and become very slow. Although Taichi-MPM is more memory efficient for high numbers of particles, NeuralMPM remains much faster, emulating 30 million particles at 25FPS.
@@ 4.3 Generalization
One notable advantage of NeuralMPM is its invariance to the number of particles, as the transition model only processes the voxelized representation. To demonstrate this, we train a model on WaterRamps, which contains about $2.3$k particles and $600$ timesteps, and evaluate it on WaterDrop-XL, which features about four times more particles, $1000$ timesteps, and no obstacles. An example snapshot is displayed in Figure 7. The larger number of particles only affects interpolation steps between the grid and particles, resulting in a negligible impact on total inference time, making the model nearly as fast despite 4 times more particles. We also validate generalization quantitatively by comparing the error rates on WaterDrop-XL of a model trained directly on it and the model trained solely on WaterRamps. With the same training budget, the latter achieves a lower MSE at $20.92\times 10^{-3}$ against $28.09\times 10^{-3}$. More trajectories are displayed in Figure 18.
@@ 4.4 Inverse design problem
Finally, we demonstrate the application of NeuralMPM for inverse problems on a toy inverse design task that consists in optimizing the direction of a ramp to make the particles reach a target location, similar to (Allen et al., 2022). We place a blob of water at different starting locations, and we then place a ramp at some location, with a random initial angle $\alpha$. The goal is to spin the ramp by tuning $\alpha$ in order to make the water end up at a desired location. The main challenges of this task are the long-range time horizon of the goal and the presence of nonlinear physical dynamics. We proceed by selecting the point where we want the water to end up and compute the average distance between the point and particles at the last simulation frame. We then minimize the distance via gradient descent, leveraging the differentiability of NeuralMPM to solve this inverse design problem. We show am example optimization in Figure 8, and additional examples in Appendix B.
@@ 5 Conclusion
We presented NeuralMPM, a neural emulation framework for particle-based simulations inspired by the hybrid Eulerian-Lagrangian Material Point Method. We have shown its effectiveness in simulating a variety of materials and interactions, its ability to generalize to larger systems and its use in inverse problems. Crucially, NeuralMPM trains in a fraction of the time it takes to train alternative approaches, and is substantially faster at inference time. By interpolating particles onto a fixed-size grid, global information is distilled into a voxelized representation that is easier to learn and process with powerful image-to-image models. The use of voxelization allows NeuralMPM to bypass expensive graph constructions, and the interpolation leads to easier generalization to a larger number of particles and constant runtime. The lack of expensive graph construction and message passing also allows for more autoregressive steps and parallel rollouts.
@@ Appendix A Training details
Our implementation, training scripts, experiment configurations, and instructions for reproducing results are publicly available at https://neuralmpm.isach.be. We implement NeuralMPM using PyTorch (Paszke et al., 2019), and use PyTorch Geometric (Fey & Lenssen, 2019) for implementing efficient particle-to-grid functions, more specifically from the Scatter and Cluster modules. For memory efficiency, we do not store all (up to) 1,000 training trajectories in memory, and rather use a buffer of about 16 trajectories over which several epochs are performed before loading a new buffer of random trajectories.
@@
The code, together with additional videos, is available at the project’s website https://neuralmpm.isach.be.
Appendix B Supplementary results
@@
In Table 2, we report the numerical MSE rollout values that were reported in the bar plots depicted in Figure 3 for Goop. Also, Figure 9 displays the error when rolling out a model for each dataset, both in terms of MSE and EMD. For both metrics, the error starts low and slowly accumulates over time. For the EMD, we use the Sinkhorn algorithm provided by (Cuturi et al., 2022).
@@
Figure 9: Error propagation during rollout. We show the mean and 80th percentile of the MSE and EMD, computed over particles and simulations, at each timestep during the rollout. The accuracy decreases as errors accumulate.
@@
Offloading gravity.
Additionally, we tried adding the gravity explicitly to the velocity output in VariableGravity, but saw degraded performance compared feeding the gravity as an input to the model for large time-bundling values.
@@
Appendix C Gallery of predicted trajectories
In addition to the trajectories in Figures 2 and 4, we show additional trajectories emulated with NeuralMPM for all datasets in Figures 12131415161718, and 19. We also release videos in the supplementary material, which we recommend watching to better see the details and limitations of NeuralMPM. This includes 10 videos per dataset of emulated trajectories on held-out test simulations.

### v3

@@ 4 Experiments
We conduct a series of experiments to demonstrate the accuracy, speed, and generalization capabilities of NeuralMPM. Specifically, we examine its robustness to hyperparameter and architectural choices through an ablation study (4.1). We compare NeuralMPM to GNS and DMCF in terms of accuracy, training time, convergence, and inference speed (4.2). We also evaluate the generalization capabilities of NeuralMPM (4.3) and illustrate how its differentiability can be leveraged to solve an inverse design problem (4.4). Through these experiments, we demonstrate that NeuralMPM is a flexible, accurate, and fast method for emulating complex particle-based simulations. The baselines established by Winchenbach & Thuerey (2024) and hybrid simulators (Li et al., 2024a; Ma et al., 2023) have promising results. However, we do not compare against them as they either use different benchmarks or are specifically tailored for certain physical domains, requiring material-specific knowledge. In contrast, NeuralMPM, like GNS and DMCF, requires only particle positions without being restricted to any particular domain. Rollouts for all experiments, models, and datasets are available at drive and in the supplementary material.
@@
Figure 3: Ablation results. Mean squared error of full rollouts on unseen test data for Goop. The default parameters are in blue. The dotted orange line ($2.4\times 10^{-3}$) indicates the MSE we obtained for GNS after $240$ hours ($20$M training steps). The dotted red line is the MSE for DMCF after the same amount of time ($5.25\times 10^{-3}$). NeuralMPM is robust to hyperparameter changes, with the biggest effects coming from the number of timesteps bundled together ($m$) and grid noise. For a rollout of length $T$, the model is called $T/m$ times, meaning lower values of $m$ require maintaining stability for longer. Autoregressive training coupled with time bundling suffices to stabilize the model, eliminating the need for noise injection. Although GNS reportedly slightly outperforms NeuralMPM, these results could not be reproduced in our experiments.
Figure 4: Example VariableGravity trajectory against baselines. Each method is unrolled starting from the initial conditions of a random test trajectory not seen during training.
@@ 4.1 Ablation study
To study the robustness of NeuralMPM to hyperparameter and architectural choices, we start with the default architecture and hyperparameters and ablate its components individually to examine their impact on performance. We vary the number K of autoregressive steps with and without grid and particle noise in the input state the number of bundled timesteps $m$ predicted by a single model call, and the depth and number of hidden channels of the network. We also investigate adding noise to stabilize rollouts, either directly to the particles’ positions or to the grid-level representation after voxelization.
Figure 3 summarizes the ablation results. A larger number $K$ of autoregressive steps yields more accurate rollouts without the need to add noise. Indeed, injecting noise does not improve accuracy and is even detrimental for $K=4$. Individually tuning the noise levels for grids and particles can modestly lower error rates, but is either very sensitive or negligible. The model performs better when bundling more timesteps, enabling faster rollouts as a single forward pass predicts more steps. We found $m=8$ to be optimal with the other default hyperparameters, outperforming larger bundling. This is because more network capacity is needed to extract information for the next $16$ or $32$ timesteps from a single state. Instead, we opted for a shallower and narrower network to balance speed and memory footprint with performance gains. In terms of network architecture, we chose a U-Net. We experiment with an FNO (Li et al., 2021) in Appendix C and find it to underperform, leading us to keep the U-Net architecture. We find the U-Net’s width and depth to have a minor impact on performance, confirming that a larger network is not needed. The grid size, however, is critical. A low resolution loses fine details, while a high resolution turns meaningful structures, such as liquid blobs or walls, into isolated voxels.
@@ 4.2 Comparison with previous work
Figure 6: Time and memory performance. Average FPS (left) and GPU VRAM usage (right) for increasing numbers of particles for a traditional solver (Taichi-MPM (Hu et al., 2018a)), NeuralMPM, and the two baselines. The two baselines quickly require very large amounts of memory and become very slow. Although Taichi-MPM is more memory efficient for high numbers of particles, NeuralMPM remains much faster, emulating 30 million particles at 25FPS. For the low particle count regime ($<10$K) we used the NeuralMPM and baselines WaterRamps models. For the high particle count regime we used untrained models and measured the throughput. The figures measures just FPS, and not the real simulation time. Taichi-MPM needs a much smaller step size than the three neural emulators ($0.2$ms vs $2.5$ms), and is therefore likely slower than all of them
@@ 4.3 Generalization
One notable advantage of NeuralMPM is that the processor is invariant to the number of particles, as the transition model only processes the voxelized representation, while both p2g and g2p scale linearly. To demonstrate this, we train a model on WaterRamps, which contains about $2.3$k particles and $600$ timesteps, and evaluate it on WaterDrop-XL, which features about four times more particles, $1000$ timesteps, and no obstacles. An example snapshot is displayed in Figure 7. The larger number of particles only affects interpolation steps between the grid and particles, resulting in a negligible impact on total inference time, making the model nearly as fast despite 4 times more particles. We also validate generalization quantitatively by comparing the error rates on WaterDrop-XL of a model trained directly on it and the model trained solely on WaterRamps. With the same training budget, the latter achieves a lower MSE at $20.92\times 10^{-3}$ against $28.09\times 10^{-3}$. More trajectories are displayed in Figure 23.
@@ 4.4 Inverse design problem
Finally, we demonstrate the application of NeuralMPM for inverse problems on a toy inverse design task that consists in optimizing the direction of a ramp to make the particles reach a target location, similar to (Allen et al., 2022). We place a blob of water at different starting locations, and we then place a ramp at some location, with a random initial angle $\alpha$. The goal is to spin the ramp by tuning $\alpha$ in order to make the water end up at a desired location. The main challenges of this task are the long-range time horizon of the goal and the presence of nonlinear physical dynamics. We proceed by selecting the point where we want the water to end up and compute the average distance between the point and particles at the last simulation frame. We then minimize the distance via gradient descent, leveraging the differentiability of NeuralMPM to solve this inverse design problem. We show an example optimization in Figure 8, and additional examples in Appendix C.
@@ 5 Conclusion
We presented NeuralMPM, a neural emulation framework for particle-based simulations inspired by the hybrid Eulerian-Lagrangian Material Point Method. We have shown its effectiveness in simulating a variety of materials and interactions, its ability to generalize to larger systems and its use in inverse problems. Crucially, NeuralMPM trains in 6% of the time it takes to train GNS and DMCF to comparable accuracy, and is 5x-10x faster at inference time. By interpolating particles onto a fixed-size grid, global information is distilled into a voxelized representation that is easier to learn and process with powerful image-to-image models. The use of voxelization allows NeuralMPM to bypass expensive graph constructions, and the interpolation leads to easier generalization to a larger number of particles and constant runtime. The lack of expensive graph construction and message passing also allows for more autoregressive steps and parallel rollouts.
@@ Appendix A Training details
Our implementation, training scripts, experiment configurations, and instructions for reproducing results are publicly available at [URL]. We implement NeuralMPM using PyTorch (Paszke et al., 2019), and use PyTorch Geometric (Fey & Lenssen, 2019) for implementing efficient particle-to-grid functions, more specifically from the Scatter and Cluster modules. For memory efficiency, we do not store all (up to) 1,000 training trajectories in memory, and rather use a buffer of about 16 trajectories over which several epochs are performed before loading a new buffer of random trajectories.
@@
The code, together with additional videos, is available at the project’s website [URL].
Appendix B Additional Discussion
Boundary conditions.
Boundary conditions are represented as stationary particles. Like moving particles, their density and zero velocity are interpolated into channels, but their positions is enforced to remain fixed during rollout. This approach allows for flexible boundary shapes, such as the ramps in WaterRamps.
Voxelization and interpolation.
The voxelization procedure used is implemented using CUDA by PyTorchCluster Fey & Lenssen (2019), p2g is implemented by us based on the voxelized representation, and g2p uses PyTorch’s bilinear interpolation (grid_sample) Paszke et al. (2019) based on the voxelized representation of the data. The cost of those operations is $\mathcal{O}(N)$, e.g., a naive implementation of voxelization assigns each point $\mathbf{p}=(x,y,z)$ to a voxel index computed as:
$\mathbf{v}=\left(\left\lfloor\frac{x}{s}\right\rfloor,\left\lfloor\frac{y}{s}\right\rfloor,\left\lfloor\frac{z}{s}\right\rfloor\right),$
[…]
Limitations of voxelization.
One of the main issues of voxelization onto a fixed regular grid means the domain is limited and it must be rectangular, like for the material point method. While a rectangular domain can be defined with local boundary conditions over the irregular domain of interest, this will be a disadvantage for sparse systems, for which Lagrangian methods might be a better fit. Likewise, out-of-bounds particles can be handled either with a larger domain, which might be more computationally expensive, by clamping the positions, or simply dropping the out-of-bounds particles. These particular conditions are known disadvantages of MPM, and of some implementations of SPH, and are therefore inherited by NeuralMPM.
Rollout stability.
Long-term stability during rollout of neural emulators is an open problem. As such, we decouple this problem from the design of the method or the architecture. There are commonly used tricks in the literature, such as noise injection, backpropagating through the autoregressive rollout, temporal bundling, having two models work in tandem (one to produce large timesteps and one to interpolate between them), using explicit constraints White et al. (2024), or modifying the gradients Schnell & Thuerey (2024). (List et al., 2025) perform a comparative study of some of these methods.
Appendix C Supplementary results
Additional results on DamBreak2D
We have also compared the accuracy and inference speed of NeuralMPM against a different implementation of GNS, and one of SEGNN, both provided by (Toshev et al., 2024), in Tables 2 and 3. As in Table 1, NeuralMPM outperforms both by a margin baselines.
[…]
Ours 20.76 2.88
GNS 114.40 224.1
SEGNN 124.39 268.4
Table 2: Full-trajectory MSE (×$\times 10^{-3}$) and Sinkhorn distance (EMD) (×$\times 10^{-4}$) for NeuralMPM, GNS, and SEGNN Brandstetter et al. (2022a) on the Dam Break 2D dataset from LagrangeBench. The two latter models are baselines provided by LagrangeBench.
Single call ($T=1$) Rollout ($T=401$)
Ours 7.41 193.50
GNS 20.46 8,170.47
SEGNN 46.04 18,194.10
Table 3: Inference time (in ms) of NeuralMPM, GNS, and SEGNN Brandstetter et al. (2022a) on the Dam Break 2D dataset from LagrangeBench. Times were averaged over all test trajectories. NeuralMPM predicts 16 frames in a single model call and still outperforms the two baselines per call, which further widens the gap for the total rollout time.
@@ Appendix C Supplementary results
In Table 4, we report the numerical MSE rollout values that were reported in the bar plots depicted in Figure 3 for Goop. Also, Figures 12 and 10 displays the error when rolling out a model for each dataset, both in terms of MSE and EMD. For both metrics, the error starts low and slowly accumulates over time. For the EMD, we use the Sinkhorn algorithm provided by (Cuturi et al., 2022).
@@
Figure 9: MSE propagation during rollout. We show the 25th, 50th and 75th percentile of the MSE, computed over particles and simulations, at each timestep during the rollout for each model and dataset. The accuracy decreases as errors accumulate. While training in 5% of the time, NeuralMPM outperform the baselines on all datasets, except WaterRamps, where it is slightly worse than GNS.
Figure 10: EMD propagation during rollout. We show the 25th, 50th and 75th percentile of the EMD, computed over particles and simulations, at each timestep during the rollout for each model dataset daset. The accuracy decreases as errors accumulate. While training in 5% of the time, NeuralMPM outperform the baselines on all datasets.
Figure 11: MSE propagation during rollout for the generalization task WaterDrop-XL. We show the 25th, 50th and 75th percentile of the MSE, computed over particles and simulations, at each timestep during the rollout for each model and daset. The accuracy decreases as errors accumulate. The models used were trained on WaterRamps and tested on WaterDrop-XL to evaluate their generalization. NeuralMPM performs better despite having a slightly worse MSE on WaterRamps than GNS.
Figure 12: EMD propagation during rollout for the generalization task WaterDrop-XL. We show the 25th, 50th and 75th percentile of the MSE, computed over particles and simulations, at each timestep during the rollout for each model and daset. The accuracy decreases as errors accumulate. The models used were trained on WaterRamps and tested on WaterDrop-XL to evaluate their generalization. NeuralMPM performs better despite having a slightly worse MSE on WaterRamps than GNS.
@@
Appendix D Gallery of predicted trajectories
We present additional rollout comparisons in Figures 15 and  16. Further, in addition to the trajectories in Figures 2 and 4, we show additional trajectories emulated with NeuralMPM for all datasets in Figures 17181920212223, and 24. We also release videos in the supplementary material, which we recommend watching to better see the details and limitations of NeuralMPM. This includes 10 videos per dataset of emulated trajectories on held-out test simulations.
Figure 15: Example MultiMaterial trajectory against baselines. Each method is unrolled starting from the initial conditions of a random test trajectory not seen during training.
Figure 16: Example WaterRamps trajectory against baselines. Each method is unrolled starting from the initial conditions of a random test trajectory not seen during training.